if 4^x+4^x-1=10 find x
Found 2 solutions by tommyt3rd, Alan3354:
Answer by tommyt3rd(5050)
Answer by Alan3354(69443)

if 4^x+4^x-1=10 find x
--------------
I think you mean
if 4^x+4^(x-1) = 10 find x
Use parentheses to make things unambiguous.
---------
if 4^x+4^x-1=10 find x
--------------
Sub y for 4^(x-1)
4y + y = 10
y = 2
-----
4^(x-1) = 2
(x-1)*log(4) = log(2)
x = 1 + log(2)/log(4)
x = 1.5